In an election, 2.8 million votes were cast and each vote was for either Candidate I or Candidate II. Candidate I received 28,000 more votes than Candidate II. What percent of the 2.8 million votes were for Candidate I?
(A) 50.05% (B) 50.1% (C) 50.5% (D) 51% (E) 55%

Answer by rfer(16322)

x+x+28000=28000000
2x=27972000
x=13986000
x+28000=14014000
------------
14014000/28000000=
14014/28000=0.5005=50.05% (A)